Which equation best models the data:
x: -2, 0, 2, 4
y: 3.5, 7, 14, 28
a)y=3.5(2)(^x/2)
b)y=3.5(2)^x
c)y=7(2)(^x/2)
d)y=7(2)^x

Answer by edjones(8007)

Let x=0, y=7
a)3.5*1 no
b)3.5*1 no
c)7*1 yes
d)7*1 yes
.
Let x= -2, y=3.5
c)7*2^-1=7*(1/2) yes
d)7*2^-2=7*(1/4) no
.
c is the answer.
.
Ed